OpenStudy (anonymous):

7 over 4- squareroot 3 simplify!

OpenStudy (earthcitizen):

\[(7/(4-\sqrt{3}))\] = \[(7/4-\sqrt{3})*(4+\sqrt{3})/(4+\sqrt{3}) = 7(4+\sqrt{3})/16-3\]

OpenStudy (earthcitizen):

\[7(4+\sqrt{3})/13\]